50,000 75o,ooo 1,000,0o0 an=0.7 100,00o 2000 2012 ak=0.3 300,000o 75,000 SIMONA 2000 soo,ooo 125,000 100,000 2012 750,000 150,000 110,000 ak - 0.5 an -0.5 Utilize the above data to compute: a) How much did capital contribute to economic growth during the time period in question for each country? b) How much did labor contribute to economic growth durin the time period in question for each country? ow much did productivity contribute to economic gro during the time period in question for each coun
d) What do these computed figures tell you about the qualitative differences in the growth experiences between these two countries? e) Explain how one would utilize Solow's growth model to analyze these qualitative differences in economic growth experiences for these countries from a theoretical perspective. f) Let's assume your above analysis transpires in a closed economy context for both countries. Explain why one of these economies would benefit more by embracing a more open economy focused adopting economic incentives designed to attract foreign direct investment in plant, equipment, infrastructure and R&D
Answer #1

First of all we need to calculate economic growth for each nation-

Milestones: Economic growth=(difference between the two years output level(Y))/current year(2012) output level.

Economic growth=250000/1000000

=0.25 (~25%)

Similarly for another nation,

Simonia: Economic growth=250000/750000

=0.33 (~33%)

(A) capital contribution-

For Milestones,

Capital contribution=difference between the two years contribution/current years contribution

=200000/300000 . =0.667 (~67%)

For Simonia,

Capital contribution=25000/150000=0.167 (~17%)

(B) labour contribution-

For Milestones,

25000/75000=0.33 (~33%)

For Simonia,

10000/110000=0.09 (~9%)

(C) contribution by productivity-

For Milestones,

250000/100000=0.25 (~25%)

For Simonia,

25000/75000=0.33 (~33%)

(D) The qualitative variation among both the economies gives us the reflection that despite a high share of capital and labor in the state of Milestones,the economic growth is more in another nation i.e. Simonia. And this difference is due to the quality based difference.